8. Let W be the subspace of R^5 defined by W={x base 1,x base 2,x base 3,

x base 4 ,x base 5) "\isin" R^5: x base 1 = 3x base 2 and x base 3 =7x base 4}. Then the basis of W is

(i) (3,1,0,0,1), (3,1,3,0,0), (3,1,0,0,1)

(ii) (3,1,0,1,1), (0,0,3,0,1), (0,0,1,3,1)

(iii) (3,1,1,0,1), (0,1,1,0,3), (0,0,1,0,1)


9. The basis of a solution space of given homogeneous linear system


X base 1 + x base 2 - x base 3 =0 X base 1 + x base 2 - x base 3 =0 X base 1 - x base 3 =0

-X base 1 + x base 3 =0 "\implies"X base 2 =0 =0 "\implies" X base 2 =0

-2X base 1 - x base 2 + 2x base 3 =0 -2X base 1 - x base 2 + 2x base 3 =0 - 2X base 1 + 2x base 3 =0

is

(i) {(1, 0, 1)}

(ii) {(1, 0, 1), (0, 1, 0)}

(iii) {(1, 1, -1), (-1, 0, 1), (-2, -1, 2)}

(iv) None

10. For a given matrix A "\begin{bmatrix}
 1 & 0 & 2 & - 3 \\
 2 & 0 & 4 & - 6 \\
 - 3 & 0 & - 6 & 9
\end{bmatrix}". Which of the following is true

(i) rank (A) =3, nullity (A) =1

(ii) rank (A) =2, nullity (A) =2

(iii) rank (A) =1, nullity (A) =3

(iv) None


1
Expert's answer
2022-02-23T14:25:32-0500

8.


"x_1=3x_2" and "x_3=7x_4"


"\begin{pmatrix}
 3x_2 \\
 x_2\\
7x_4\\
x_4\\
x_5
\end{pmatrix}=x_2\begin{pmatrix}
 3 \\
 1 \\
0\\
0\\
0
\end{pmatrix}+x_4\begin{pmatrix}
 0\\
 0 \\
7\\
1\\
0
\end{pmatrix}+x_5\begin{pmatrix}
 0 \\
 0\\
0\\
0\\
1
\end{pmatrix}"


Basis of W is "(3,1,0,0,0),(0,0,7,1,0),(0,0,0,0,1)"



9.

"x_1+x_2-x_3=0.....(i)\\
x_1-x_3=0......(ii)\\
-2x_1-x_2+2x_3=0....(iii)\\
-2x_1+2x_3=0....(iv)"



"-2(ii)=(iv)," therefore ignoring "(iv)"



"\begin{pmatrix}
 1&&1&&-1 \\
 1&&0&&-1\\
-2&&-1&&2
\end{pmatrix}\begin{pmatrix}
 x_1 \\
 x_2 \\
x_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
0
\end{pmatrix}"



"\implies\begin{pmatrix}
 1&&0&&-1 \\
 0&&1&&0\\
0&&0&&0
\end{pmatrix}\begin{pmatrix}
 x_1 \\
 x_2\\
x_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
0
\end{pmatrix}"



"x_1=x_3\\
x_2=0"


"\begin{pmatrix}
 x_1 \\
 0 \\
x_1
\end{pmatrix}=x_1\begin{pmatrix}
 1 \\
 0\\
1
\end{pmatrix}"


Basis is "[1,0,1]"




10.


"\begin{pmatrix}
 1&&0&&2&& -3 \\
 2&&0&&4&&-6 \\
-3&&0&&-6&&9
\end{pmatrix}"



"\frac{1}{2}R_2\to\>R_2"


"\frac{-1}{3}R_3\to\>R_3"


"\begin{pmatrix}
 1&&0&& 2&&-3 \\
 1&&0&&2& & -3\\
1&&0&&2&&-3
\end{pmatrix}"


"R_2-R_1\to\>R_2\\
R_3-R_1\to\>R_3"



"\begin{pmatrix}
 1&&0&&2&&-3 \\
 0&&0& & 0&&0\\
0&&0&&0&&0
\end{pmatrix}"


Rank "=1"

Nullity "=3"